# Tiered Reward Functions: Specifying and Fast Learning of Desired Behavior

### v2

@@ 7 Experiments
In this section, we aim to verify the usefulness of our proposed Tiered Reward. We empirically show that Tiered Rewards are able to make learning faster on multiple tabular domains. We extend our results to environments with high-dimensional inputs and show that Tiered Reward works with deep RL algorithms. Finally, we also explore the influence of the number of tiers and find that having more tiers can induce faster learning.
Fast Learning with Tiered Rewards
Guaranteeing Pareto-optimal behavior is not the sole benefit of using Tiered Rewards; we find that it also leads to fast learning. There are many ways to design a Tiered Reward because it is a class of reward functions that is only constrained by an inequality (Definition 6.1), and not by specific reward values. In this section, for simplicity and clarity we use $k$-Tiered Reward defined by:
@@
Figure 2: Left: A Tiered Reward in a grid world with $6$ tiers. Start state is in the bottom left corner, goal state in top right. Darker colors correspond to more negative reward values. Right: Flag grid by Ng, Harada, and Russell (1999).
@@
The “Flag Grid” from Ng, Harada, and Russell (1999) is a natural $6$-Tier MDP to study Tiered Reward. In this grid world (Figure 2 right), the agent begins in the bottom-left corner and must learn to pick up four flags in sequence before reaching the goal. The state space is the location of the agent plus the flag it has collected. The agent can move in $4$ directions with an $80\%$ success rate, while acting randomly $20\%$ of the time. All states in which the agent possesses the same number of flags constitute a tier, totaling 5 tiers. The goal constitutes the sixth and final tier.
To evaluate Tiered Reward, we compare it against two baseline reward functions. The first one, following Koenig and Simmons (1993), we call action penalty. This reward penalizes each step with $-1$, until the goal state is reached and the agent is awarded $+1$. To reiterate, such step-wise negative reinforcement encourages directed exploration. This reward only assumes knowledge about the position of the goal.
The Flag Grid domain was introduced in the context of potential-based shaping (Ng, Harada, and Russell 1999). In that work, the authors showed how knowledge of the subgoals could be leveraged through shaping rewards to guide the learning process. Our Tiered Rewards can be used similarly, so we compare potential-based shaping to Tiered Rewards in this environment. For direct comparability, we use the Tiered Reward as a potential function $\Phi(s)=R_{tier}(s)$ to shape the action-penalty reward, resulting in what we call tier-based shaping reward: $R_{tbs}(s,a,s^{\prime})=R_{penalty}(s)+\gamma\Phi(s^{\prime})-\Phi(s)$.
@@
Figure 3: Q Learning curves for Flag Grid. The two plots use different initial Q-values. Each experiment was run with $30$ seeds. We experimented with a range of discounts $\gamma\in\{0.99,0.95,0.90,0.85\}$ and found they all had similar performances, and only report $\gamma=0.90$ here.
@@
We compare these reward functions first in the context of Q-Learning (Watkins and Dayan 1992), arguably the most well-understood and widely applicable RL algorithm. As Figure 3 (top) shows, Tiered Reward always learns fastest of the three. Perhaps surprisingly, tier-based shaping reward performs orders of magnitudes worse than even action penalty. That is because, with discounting, the shaping function $F(s,a,s^{\prime})=\gamma\Phi(s^{\prime})-\Phi(s)$ becomes positive when $s$ and $s^{\prime}$ belong to the same tier. As a result, the zero-initialized Q-values are not optimistic with respect to these rewards, and so exploration with $R_{tbs}$ is undirected, leading to slow learning. For a fairer comparison, we also plot the learning curves where all Q-values are initialized to some arbitrary large value so that $R_{tbs}$ also enjoys directed exploration as the two other rewards (Figure 3, bottom). We tried different initialization values ($10^{10},10^{5},10^{3}$) and observed similar results, so we only report $Q_{init}=10^{5}$ here. There is a noticeable spike in time taken to reach the goal early on during training. It is the result of optimistic Q-value initialization, which leads to more exploration and thus slower learning. Regardless of this trade off, Tiered Reward still consistently outperforms the two baselines.
@@
It is important to note that our goal here is not to argue shaping is ineffective, nor to determine how to initialize Q-values for fast learning, but solely to demonstrate the usefulness of Tiered Reward in various different settings. To start, it makes learning faster than tier-based shaping reward and action penalty for different discount factors and Q-value initialization schemes. Moreover, it is simple to design and implement; there is no need to engineer environment-specific reward and initial Q-values to accelerate learning.
Tiered Reward in Deep RL
@@
Figure 4: Illustration of EmptyGrid, FourRooms, and DoorKey from left to right. The agent is in red, and the goal in green. All three environments use visual observations.
We further show that Tiered Reward also make learning faster in Deep RL with image observations. We choose three goal—obstacle environments from MiniGrid (Chevalier-Boisvert et al. 2023):
  1. 1.
    EmptyGrid: a visual grid world ($8\times 8$) where the agent starts in one corner and aims for the opposite corner.
  2. 2.
    FourRooms: a visual grid world ($19\times 19$) where the agent must navigate to the goal (green) in one of four rooms by going through the gaps in the wall.
  3. 3.
    DoorKey: a visual grid world ($8\times 8$) where the agent must first pick up a key, then travel to and unlock a door, and finally navigate to the goal.
In all three environments (Figure 4), the agent aims to learn a policy to navigate to the goal using image observations. FourRooms is a long-horizon problem; the DoorKey environment has a complicated transition function and action space, and is hard to solve using classical RL algorithms with a sparse reward. In all three environments, we design Tiered Reward with three tiers. For EmptyGrid and FourRooms, tiers are assigned based on each state’s $L1$ distance to the goal; for DoorKey, tiers are assigned based on the sub-goals the agent has completed (getting key, opening door, and reaching goal).
We use Proximal Policy Optimization (Schulman et al. 2017) (PPO) as the RL algorithm, and use an ImpalaCNN (Espeholt et al. 2018) as the policy network architecture. More experiment details are included in Appendix E. To provide numerical stability, deep RL methods often employ reward scaling and clipping (Henderson et al. 2018). To follow, we linearly scale the tiered reward values to be between $-1$ and 00.
Following the previous set up, we compare Tiered Reward against action penalty and tier-based shaping. To evaluate the learning performance, we plot the episodic return during learning using the original reward function from MiniGrid (Chevalier-Boisvert et al. 2023). The original reward functions in MiniGrid are designed by human experts for each MiniGrid environment, and express the task of reaching the goal quickly and avoiding obstacles. By plotting the learned policy’s performance on the original reward function, we can examine how quickly the agent has learned the desired behavior. Figure 5 shows the learning performance of PPO on three environments. In all three, Tiered Reward is able to accumulate high episodic returns, thereby learning the desired behavior. Across the three environments, Tiered Reward show the fastest learning, outperforming the Action Penalty and Tier Based Shaping baselines. Tier Based Shaping performs differently according to how hard the exploration problem is: as discussed before, when $s$ and $s^{\prime}$ are in the same tier, the shaping function $F(s,a,s^{\prime})=\gamma\Phi(s^{\prime})-\Phi(s)>0$, and therefore encourages the agent to exploit rather than explore. For FourRooms, Tiered Reward initially learns slower than Tier Based Shaping, but later catches on and eventually outperforms its counterpart. All three reward functions suffer from large standard deviation in this environment likely because this environment is a hard exploration problem; different exploration during learning leads to wildly different outcomes.
@@
Influence of More Tiers
Figure 6: Left: Frozen lake environment with holes in dark blue. Right: Wall Grid environment with walls in black and lava in orange.
Figure 7: Q-Learning results with $\gamma=0.9$. Each experiment was run with $300$ seeds.
@@
    Frozen Lake: a slippery grid with holes that will swallow the agent (Figure 6 left). The objective is to reach the goal without falling into any holes. Each of the $4$ directional actions succeed $1/3$ of the time, and slip to either side with probability $1/3$.
@@
    Wall Grid: a grid world with multiple lava states and wall states (Figure 6 right). The agent has to circle around the walls while avoiding the lava to get to the goal. Transition dynamics same as Grid.
For Chain and Grid, tiers are decided based on their $L1$ distance to the goal; for Frozen Lake and Wall Grid, tiers are assigned based on the sum of $L1$ distance to the goal and start state, weighted $2:1$. Since there are $3$ types of states in the latter two environments, we start at $3$ tiers instead of $2$.
In absence of a “correct” reward function that expresses the tasks, we measure the learning speed by recording the steps required for the agent to reach the goal for the first time. For fair comparison, we optimistically initialize $Q_{init}=10^{5}$ so that $R_{tbs}$ also enjoys the benefit of directed exploration. The results are presented in Figure 7. We repeat the same experiments with a model-based RL algorithm, RMAX (Brafman and Tennenholtz 2002). The results are similar to that of Q-learning (Figure 11 in Appendix F). As expected, Tiered Reward makes learning faster as the number of tiers increases because more information about the environment aids reward design. Tiered Reward consistently beats action penalty and is at least as good as tier-based shaping reward, and often much better. Even when Tiered Reward performs the same as shaped reward, it provides the added benefit of simplicity and better interpretability—it is based only on states, and not state–action–state triples.
@@
Adding more tiers is more complicated in deep RL settings because scaling the reward values to $[-1,0]$ for learning stability in deep RL will decrease the difference of reward value among tiers, especially among higher tiers. Figure 8 shows the learning performance of Tiered Reward with $3$, $5$, $7$, and $9$ tiers on EmptyGrid, FourRooms, and DoorKey. In EmptyGrid, having more tiers makes learning faster and more stable. In DoorKey, $7$ and $9$ tiers produce similar results, both significantly faster than $3$ or $5$ tiers. In FourRooms, $5$ tier leads to the fastest learning result, and using more than $5$ tiers monotonically decreases learning performance. We believe this is because of numerical issues during scaling of Tiered Reward: reward values for all states in tiers higher than $3$ gets normalized to very close to 00 (in the range $10^{-6}$ to $10^{-15}$), and therefore tiers become indistinguishable under some numerical accuracy. Having more tiers suffers more from this problem because more states become indistinguishable [^1] To alleviate the numerical issues caused by scaling, we set discount factor to $0.5$ in these experiments to obtain smaller original Tiered Reward. See Appendix H for further details. We hypothesize that this issue can be resolved with a smarter reward scaling or clipping method, and leave that for future work. In practice, the optimal number of tiers can be determined empirically or a priori with MDP-specific information. Though more tiers sometimes leads to slower learning than $3$ tiers, they all provide faster learning than the two reward baselines in Figure 5.
@@ 8 Conclusion
In contrast to standard reward-design solutions that are environment-dependent, we presented Tiered Rewards—a class of environment-independent reward functions that provably leads to (Pareto) optimal behavior and empirically leads to fast learning. Tiered Reward can be defined in both tabular and high dimensional environments, and is RL-algorithm agnostic. One interesting future work is getting theoretical guarantees that Tiered Reward lead to asymptotically faster learning.
@@
$\begin{split}\sum_{t=0}^{\infty}(g_{t}-g_{t}^{*})f_{t}^{g}&=\sum_{t=0}^{\infty}\sum_{j=0}^{t}(g_{j}-g_{j}^{*})f_{t}^{g}-\sum_{t=0}^{\infty}\sum_{j=0}^{t-1}(g_{j}-g_{j}^{*})f_{t}^{g}\\ &=\sum_{t=0}^{\infty}\sum_{j=0}^{t}(g_{j}-g_{j}^{*})f_{t}^{g}-\sum_{t=1}^{\infty}\sum_{j=0}^{t-1}(g_{j}-g_{j}^{*})f_{t}^{g}\\ &=\sum_{t=0}^{\infty}\sum_{j=0}^{t}(g_{j}-g_{j}^{*})f_{t}^{g}-\sum_{t’=0}^{\infty}\sum_{j=0}^{t’}(g_{j}-g_{j}^{*})f_{t’+1}^{g}\\ &=\sum_{t=0}^{\infty}(\sum_{j=0}^{t}g_{j}-g_{j}^{*})(f_{t}^{g}-f_{t+1}^{g})\\ \end{split}$
@@
Note that the (**)(**)( * * ) step is justified only because $\sum_{j=0}^{t}p_{j}^{m}-{p_{j}^{*}}^{m}\geq 0,\forall m=[2..k],\forall t$. The inequalities show that $\pi$ achieves higher reward than the optimal policy, which is a contradiction. No such $\pi$ exists. ∎
@@ Appendix C Visualization of Pareto-optimal Policies
We visualize each policy as a shaded area upper bounded by $\sum_{t}g_{t}$ and lower bounded by $-1+\sum_{t}o_{t}$ in Figure 9. This visualization can be understood as separating the probability space into two, with the goal-reaching probability on the top half of the y-axis in $[0,1]$ and obstacle-hitting probability in the bottom half of the y-axis in $[-1,0]$. With this visualization, a Pareto-dominated policy will cover an area that is entirely enclosed by that of a dominating policy because of lower goal-reaching probabilities on the top half and higher obstacle-hitting probabilities on the bottom half. As Figure 9 shows, “always right” and “always left” do not cover each other, so they are incomparable. Specifically, “always right” has a slightly higher probability of reaching the goal (brown), but “always left’ has a lower probability of reaching the lava (purple and teal).
For comparison, we plot another policy, which we call $R$, that is state-dependent and moves in the direction of the goal. For policy $R$, the probability of reaching the target increases with time because each step has a 20% slip probability; agents could slip early on and take longer to reach the goal. Note that area covered by $R$ (red, brown, orange, and purple) completely subsumes that of “always right” (brown and orange), demonstrating that “always right” is dominated by $R$. “Always left”, on the other hand, is not dominated by $R$ because it has a lower probability of reaching lava (teal). However, “always left” is not Pareto-optimal of course, because it is dominated by policy $G$ in Table 1.

### v3

@@ 6 Experiments
Guaranteeing Pareto-optimal behavior is not the sole benefit of using Tiered Rewards; we find that it also leads to fast learning compared to two baseline reward functions. The first one, following Koenig & Simmons (1993), we call action penalty. This reward penalizes each step with $-1$, until the goal state is reached and the agent is awarded $+1$. To reiterate, such step-wise negative reinforcement encourages directed exploration assuming only knowledge about position of the goal. The second one uses reward shaping as  Ng et al. (1999) showed that subgoals can be leveraged through shaping rewards to guide the learning process. Reward shaping using an optimal value function is not a fair comparison both because it contains more information about the environment than tiers and it requires solving the environment with a pre-specified ground truth reward function, which induces a circular logic in reward design. For direct comparability, we use the Tiered Reward as a potential function $\Phi(s)=R_{tier}(s)$ to shape the action-penalty reward, resulting in what we call tier-based shaping reward: $R_{tbs}(s,a,s^{\prime})=R_{penalty}(s)+\gamma\Phi(s^{\prime})-\Phi(s)$.
There are many ways to design a Tiered Reward because it is a class of reward functions that is only constrained by an inequality (Definition 5.4), and not by specific reward values. In this section, for simplicity and clarity we use $k$-Tiered Reward defined by:
@@
Empirically, we show Tiered Reward provides faster learning on multiple tabular domains. We further extend our results to environments with high-dimensional image inputs and deep RL algorithms. Finally, we also explore the influence of the number of tiers and find that having more tiers can induce faster learning.
6.1 Fast Learning with Tiered Reward
@@
The “Flag Grid” from Ng et al. (1999) is a natural $6$-Tier MDP to study Tiered Reward. In this grid world (Figure 2(a)), the agent begins in the bottom-left corner and must learn to pick up four flags in sequence before reaching the goal. The agent can move in $4$ directions with an $80\%$ success rate, while acting randomly $20\%$ of the time. All states in which the agent possesses the same number of flags constitute a tier, totaling 5 tiers. The goal constitutes the sixth and final tier.
(a) Flag Grid. The state space is the location of the agent plus the flag it has collected.
@@
(b) Q Learning curves for Flag Grid. The two plots use different initial Q-values. $\gamma\in\{0.99,0.95,0.90,0.85\}$ and $Q_{init}\in\{10^{10},10^{5},10^{3}\}$ all had similar performances, so we only report $\gamma=0.90,Q_{init}=10^{5}$ here. Error bar show standard deviation from $30$ seeds.
Figure 2: Tiered Reward leads to fast learning on the Flag Grid.
@@
We evaluate all three reward functions with Q-Learning (Watkins & Dayan, 1992), arguably the most well-understood and widely applicable RL algorithm. As Figure 2(b) (left) shows, Tiered Reward learns fastest of the three. Perhaps surprisingly, Tier Based Shaping performs orders of magnitudes worse than even action penalty. That is because, with discounting, the shaping function $F(s,a,s^{\prime})=\gamma\Phi(s^{\prime})-\Phi(s)$ becomes positive when $s$ and $s^{\prime}$ belong to the same tier. As a result, zero-initialized Q-values are not optimistic with respect to these rewards, and so exploration with $R_{tbs}$ is undirected and slow. For a fairer comparison, we also plot the learning curves where all Q-values are initialized to some arbitrary large value so that $R_{tbs}$ also enjoys directed exploration as the two other rewards (Figure 2(b), right). There is a noticeable spike in time taken to reach the goal early on during training. It is the result of optimistic Q-value initialization, which leads to more exploration and thus slower learning. Regardless of this trade off, Tiered Reward still consistently outperforms the two baselines.
@@
It is important to note that our goal here is not to argue shaping is ineffective, nor to determine how to initialize Q-values for fast learning, but solely to demonstrate the usefulness of Tiered Reward in various different settings. To start, it makes learning faster than tier-based shaping reward and action penalty for different discount factors and Q-value initialization schemes. Moreover, it is simple to design and implement; there is no need to engineer environment-specific reward structures and initial Q-values to accelerate learning.
6.2 Tiered Reward in Deep RL
@@
Figure 3: Illustration of EmptyGrid, FourRooms, and DoorKey from left to right. The agent is in red, and the goal in green. The objective is to navigate to the goal; in FourRooms, the agent has to find the gaps in the wall; in DoorKey, the agent must first pick up the key and unlock the yellow door. All three environments use visual observations.
@@
We further show that Tiered Reward also make learning faster in Deep RL with image observations. We choose three goal–obstacle environments from MiniGrid (Chevalier-Boisvert et al., 2023): EmptyGrid, FourRooms, and DoorKey. In all three environments (Figure 3), the agent aims to learn a policy to navigate to the goal using image observations. FourRooms is a long-horizon problem; the DoorKey environment has a complicated transition function and action space, and is hard to solve using classical RL algorithms with a sparse reward. In all three environments, we design Tiered Reward with three tiers. For EmptyGrid and FourRooms, tiers are assigned based on each state’s $L1$ distance to the goal; for DoorKey, tiers are assigned based on the sub-goals the agent has completed (getting key, opening door, and reaching goal).
We use Proximal Policy Optimization (Schulman et al., 2017) (PPO) as the RL algorithm. To provide numerical stability, deep RL methods often employ reward scaling and clipping (Henderson et al., 2018). To follow, we linearly scale the tiered reward values to be between $-1$ and 00. More experiment details are included in Appendix F.
For evaluation, we plot the episodic return with respect to the original reward function from MiniGrid (Chevalier-Boisvert et al., 2023). The original reward functions in MiniGrid are designed by human experts and express the task of reaching the goal quickly and avoiding obstacles. Thus, performance on the original reward function shows how quickly the agent has learned the specified behavior. Figure 4 shows the learning performance of PPO trained on different reward functions on MiniGrid. Across the three environments, Tiered Reward show the fastest learning, outperforming the Action Penalty and Tier Based Shaping baselines. Tier Based Shaping performs differently according to how hard the exploration problem is: as discussed before, when $s$ and $s^{\prime}$ are in the same tier, the shaping function $F(s,a,s^{\prime})=\gamma\Phi(s^{\prime})-\Phi(s)>0$, and therefore encourages the agent to exploit rather than explore. For FourRooms, Tiered Reward initially learns slower than Tier Based Shaping, but later catches on and eventually outperforms its counterpart. All three reward functions suffer from large standard deviation in FourRooms likely because this environment is a hard exploration problem; different exploration during learning leads to wildly different outcomes.
6.3 Influence of More Tiers
(a) Top: Frozen lake environment with holes in dark blue.
Bottom: Wall Grid environment with walls in black and lava in orange.
(b) Q-Learning results with $\gamma=0.9$. Since there are $3$ types of states in the latter two environments, we start at $3$ tiers instead of $2$. Error bar show standard deviation from $300$ seeds.
Figure 5: Influence of more tiers on four grid worlds with Q learning. RMAX results in Figure 9.
@@ 6.3 Influence of More Tiers
    Frozen Lake: a slippery grid with holes that will swallow the agent (Figure 5(a) top). The objective is to reach the goal without falling into any holes. Each of the $4$ directional actions succeed $1/3$ of the time, and slip to either side with probability $1/3$.
@@
    Wall Grid: a grid world with multiple lava states and wall states (Figure 5(a) bottom). The agent has to circle around the walls while avoiding the lava to get to the goal. Transition dynamics same as Grid.
For Chain and Grid, tiers are decided based on their $L1$ distance to the goal; for Frozen Lake and Wall Grid, tiers are based on the sum of $L1$ distance to the goal and start state, weighted $2:1$.
In absence of a “correct" reward function that expresses the tasks, we measure the learning speed by recording the steps required for the agent to reach the goal for the first time. For fair comparison, we optimistically initialize $Q_{init}=10^{5}$ so that $R_{tbs}$ also enjoys the benefit of directed exploration. The results are presented in Figure 5(b). We repeat the same experiments with a model-based RL algorithm, RMAX (Brafman & Tennenholtz, 2002). The results are similar to that of Q-learning (Figure 9 in Appendix E). As expected, Tiered Reward makes learning faster as the number of tiers increases because more information about the environment aids reward design. Tiered Reward consistently beats action penalty and is at least as good as tier-based shaping reward, and often much better. Even when Tiered Reward performs the same as shaped reward, it provides the added benefit of simplicity and better interpretability—it is based only on states, and not $(s,a,s^{\prime})$ triples.
@@
Adding more tiers achieves similar effects in deep RL settings under numerical constraints. Figure 6 shows the learning performance of Tiered Reward with $3$, $5$, $7$, and $9$ tiers on MiniGrid environments. In EmptyGrid, having more tiers makes learning faster and more stable. In DoorKey, $7$ and $9$ tiers produce similar results, both significantly faster than $3$ or $5$ tiers. In FourRooms, $5$ tier leads to the fastest learning result, and using more than $5$ tiers monotonically decreases learning performance. We believe this is because of numerical issues during scaling of Tiered Reward: we scale reward values to $[-1,0]$ for stability in learning, but it also decreases the difference of reward value among tiers, especially among higher (more desirable) tiers. Reward values for all states in tiers higher than $3$ gets normalized to very close to 00 (in the range $10^{-6}$ to $10^{-15}$), and therefore tiers become indistinguishable under some numerical accuracy. Having more tiers suffers more from this problem because more states become indistinguishable, and as a result have worse performance.  [^1] To alleviate the numerical issues caused by scaling, we set discount factor to $0.5$ in these experiments to obtain smaller original Tiered Reward. See Appendix H for further details We hypothesize that this issue can be resolved with a smarter reward scaling or clipping method, and leave that for future work. In practice, the optimal number of tiers can be determined a priori with MDP-specific information or empirically. Though more tiers sometimes leads to slower learning than $3$ tiers, all Tiered Rewards provide faster learning than the two reward baselines in Figure 4.
@@ 7 Conclusion
In contrast to standard reward-design solutions that are environment-dependent, we presented Tiered Rewards—a class of environment-independent reward structures that provably leads to (Pareto) optimal behavior and empirically leads to fast learning. Tiered Reward can be defined in both tabular and high dimensional environments, and is RL-algorithm agnostic. Interesting future work includes getting theoretical guarantees that Tiered Reward lead to asymptotically faster learning and addressing the numerical issues with more tiers.
Acknowledgments
This research is supported in part by the Office of Naval Research (ONR) award N00014-20-1-2115.
@@ Acknowledgments
$\begin{split}\sum_{t=0}^{\infty}(g_{t}-g_{t}^{*})f_{t}^{g}&=\sum_{t=0}^{\infty}\sum_{j=0}^{t}(g_{j}-g_{j}^{*})f_{t}^{g}-\sum_{t=0}^{\infty}\sum_{j=0}^{t-1}(g_{j}-g_{j}^{*})f_{t}^{g}\\ &=\sum_{t=0}^{\infty}\sum_{j=0}^{t}(g_{j}-g_{j}^{*})f_{t}^{g}-\sum_{t=1}^{\infty}\sum_{j=0}^{t-1}(g_{j}-g_{j}^{*})f_{t}^{g}\\ &=\sum_{t=0}^{\infty}\sum_{j=0}^{t}(g_{j}-g_{j}^{*})f_{t}^{g}-\sum_{t\textquoteright=0}^{\infty}\sum_{j=0}^{t\textquoteright}(g_{j}-g_{j}^{*})f_{t\textquoteright+1}^{g}\\ &=\sum_{t=0}^{\infty}(\sum_{j=0}^{t}g_{j}-g_{j}^{*})(f_{t}^{g}-f_{t+1}^{g})\\ \end{split}$
@@
Note that the ()(**)( ∗ ∗ ) step is justified only because $\sum_{j=0}^{t}p_{j}^{m}-{p_{j}^{*}}^{m}\geq 0,\forall m=[2..k],\forall t$. The inequalities show that $\pi$ achieves higher reward than the optimal policy, which is a contradiction. No such $\pi$ exists. ∎
@@ Appendix C Visualization of Pareto-optimal Policies
We visualize each policy as a shaded area upper bounded by $\sum_{t}g_{t}$ and lower bounded by $-1+\sum_{t}o_{t}$ in Figure 7. This visualization can be understood as separating the probability space into two, with the goal-reaching probability on the top half of the y-axis in $[0,1]$ and obstacle-hitting probability in the bottom half of the y-axis in $[-1,0]$. With this visualization, a Pareto-dominated policy will cover an area that is entirely enclosed by that of a dominating policy because of lower goal-reaching probabilities on the top half and higher obstacle-hitting probabilities on the bottom half. As Figure 7 shows, “always right" and “always left" do not cover each other, so they are incomparable. Specifically, “always right" has a slightly higher probability of reaching the goal (brown), but “always left’ has a lower probability of reaching the lava (purple and teal).
For comparison, we plot another policy, which we call $R$, that is state-dependent and moves in the direction of the goal. For policy $R$, the probability of reaching the target increases with time because each step has a 20% slip probability; agents could slip early on and take longer to reach the goal. Note that area covered by $R$ (red, brown, orange, and purple) completely subsumes that of “always right” (brown and orange), demonstrating that “always right" is dominated by $R$. “Always left”, on the other hand, is not dominated by $R$ because it has a lower probability of reaching lava (teal). However, “always left” is not Pareto-optimal of course, because it is dominated by policy $G$ in Table 1.